OpenStudy (anonymous):

how do your graph this y<=-2x-5 y>5x+3

OpenStudy (danjs):

The way i would do it is find the two axis intercept points

OpenStudy (danjs):

Let x = 0, calculate y for the first point (o,y)

OpenStudy (danjs):

THen Let Y = 0 and find x, that is the second point (x,0) 2 points define a line

OpenStudy (danjs):

For example \[y \le -2x - 5\] Let x = 0

OpenStudy (danjs):

y = 0, x = -5/2 , point (-5/2, 0) is on the line, so graphing,

OpenStudy (danjs):



OpenStudy (danjs):

Solid line for Less than or Equal to Dotted Line for less than or greater than

OpenStudy (anonymous):

thanks can you help me on more to

OpenStudy (danjs):

sure

OpenStudy (danjs):

pose up a new question

OpenStudy (anonymous):

wait is that a -5/20

OpenStudy (danjs):

no the point x=-5/2 y=0, the x-intercept

OpenStudy (danjs):

Let Y=0, solve for x, you get -5/2

OpenStudy (danjs):

Let x = 0, solve for Y, you get y=-5

OpenStudy (danjs):

2 points

OpenStudy (anonymous):

o ok ill delete this one and post a new one

OpenStudy (danjs):

just close it i think